he annual per capita consumption of bottled water was 31.9 gallons. Assume that the per capita consumption of bottled water is approximately normally distributed with a mean of 31.9 and a standard deviation of 10 gallons.
a. What is the probability that someone consumed more than 37 gallons of bottled​ water?
b. What is the probability that someone consumed between 20 and 30 gallons of bottled​ water?
c. What is the probability that someone consumed less than 20 gallons of bottled​ water?
d. 99​% of people consumed less than how many gallons of bottled​ water?
1
Expert's answer
2020-12-01T20:45:34-0500

"(a) \; P(x > 37) = 1 – P(x < 37) \\

= 1 – P(\frac{x – μ}{σ} < \frac{37 – 31.9}{10}) \\

= 1 – P(z < 0.51) \\

= 0.3050"

"(b)\; P( 20 < x < 30) = P(\frac{20 – 31.9}{10} < \frac{x – μ}{σ} < \frac{30 – 31.9}{10}) \\

= P(-1.19 < z < -0.19) \\

= P(z < -0.19) – P(z < -1.19) \\

= 0.4246 – 0.1170 \\

= 0.3076"

"(c) \; P(x < 20) = P(\frac{x – μ}{σ} < \frac{20 – 31.9}{10}) \\

= P(z < -1.19) \\

= 0.1170 \\

(d) \; P(z < 2.33) = 0.99\\

z = 2.33 \\

x = z \times σ + μ \\

x = 2.33 \times 10 + 31.9 = 55.2"

Answer: 55.2 gallons